HETAV PATEL

Grade 11,

Angular momentum in second bohr orbit of H -atom is x . Then find out angular momentum in 1st excited state of li^+2 ion

Angular momentum in second bohr orbit of H -atom is x . Then find out angular momentum in 1st excited state of li^+2 ion
 

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
Angular momentum of 2nd Bohr orbit is x.
So, L = nh/2π = 2h/2π = h/π
For, Li2+, it is 2st excited state. This also means that n = 2.
As angular momentum does not depend on atomic number,
So, L = nh/2π = 2h/2π = h/π
Thus angular momentum of both the systems are same, i.e. x
The answer is x.